# AdaRC: Mitigating Graph Structure Shifts during Test-Time

### v1

@@ 5 $\to$ 2 10 $\to$ 2 20 $\to$ 2 2 $\to$ 20 5 $\to$ 20 10 $\to$ 20
$+$ AdaRC 88.78 $\pm$ 0.13 88.55 $\pm$ 0.44 88.10 $\pm$ 0.21 97.01 $\pm$ 1.00 97.24 $\pm$ 1.11 97.89 $\pm$ 0.25
C.4 Robustness to additional adversarial shift
While AdaRC primarily targets natural structure shifts, inspired by [14], we test the robustness of AdaRC against adversarial attacks by applying the PR-BCD attack [10] on the target graph in our Syn-Cora experiments, varying the perturbation rate from 5% to 20%. The results are shown in Table 5. We found that while the accuracy of ERM dropped by 20.2%, the performance of AdaRC only decreased by 2.3%. This suggests that our algorithm has some robustness to adversarial attacks, possibly due to the overlap between adversarial attacks and structure shifts. Specifically, we observed a decrease in homophily in the target graph under adversarial attack, indicating a similarity to structure shifts.
@@ C.4 Robustness to additional adversarial shift
C.5 Ablation study with different loss functions
We compare our proposed PIC loss with two existing surrogate losses: entropy [33] and pseudo-label [22]. While PIC loss use the ratio form of $\sigma_{\text{intra}}^{2}$ and $\sigma_{\text{inter}}^{2}$, we also compare it with a difference form $\sigma_{\text{intra}}^{2}-\sigma_{\text{inter}}^{2}$, which also encourage larger $\sigma_{\text{inter}}^{2}$ and smaller $\sigma_{\text{intra}}^{2}$. The results are shown in Table 6: Our PIC loss has better performance under four structure shift scenarios.
@@ C.5 Ablation study with different loss functions
C.6 Hyperparameter sensitivity with different number of GPR steps $K$
Although the AdaRC does not involve any hyperparameters other than the learning rate $\eta$ and number of adaptation rounds $T$, it may be combined with GNN models with different dimension of ${\bm{\gamma}}$. Therefore in this part, we combine AdaRC with GPRGNN models using different $K$, i.e., the number of GPR steps, to test the robustness to different hyperparameter selection of the GNN model. Specifically, we tried values of $K$ ranging from 3 to 15 on Syn-Cora and Syn-Products datasets. Notice that in our experiments in 5.1, we use $K=9$. As shown in Table 7, AdaRC remains effective under a wide range of $K$.
@@ C.6 Hyperparameter sensitivity with different number of GPR steps $K$
$+$ AdaRC 72.09 $\pm$ 0.50 71.42 $\pm$ 0.65 70.58 $\pm$ 1.01 69.69 $\pm$ 1.06 69.48 $\pm$ 1.16 69.35 $\pm$ 0.66 69.72 $\pm$ 0.70
C.7 Computation time
Due to the need to adapt hop-aggregation parameters ${\bm{\gamma}}$, AdaRC inevitably introduces additional computation costs, which vary depending on the chosen model, target graph, and base TTA algorithm. We documented the computation times for each component of ERM + AdaRC and T3A + AdaRC in our CSBM experiments:
@@ C.7 Computation time
    • Forward pass involves calculation of ${\bm{Z}}$ using the current ${\bm{\gamma}}$ and cached $\{{\bm{H}}^{(0)},\cdots,{\bm{H}}^{(K)}\}$, and prediction using the linear layer classifier (or with T3A algorithm). Since AdaRC only updates ${\bm{\gamma}}$, $\{{\bm{H}}^{(0)},\cdots,{\bm{H}}^{(K)}\}$ can be cached without recomputation in each epoch. Note that other TTA algorithms could also adopt the same or similar caching strategies.
@@
ERM $+$ AdaRC Adaptation (for each epoch) 3.292 $\pm$ 0.254 11.89%
- Forward pass 1.224 $\pm$ 0.131 4.42%
- Computing PIC loss 0.765 $\pm$ 0.019 2.76%
- Back-propagation 1.189 $\pm$ 0.131 4.30%
- Updating parameter 0.113 $\pm$ 0.001 0.41%
T3A $+$ AdaRC Adaptation (for each epoch) 6.496 $\pm$ 0.333 23.46%
- Forward pass 4.464 $\pm$ 0.248 16.12%
- Computing PIC loss 0.743 $\pm$ 0.011 2.68%
- Back-propagation 1.174 $\pm$ 0.167 4.24%
- Updating parameter 0.115 $\pm$ 0.004 0.41%
We provide the computation time for each stage in Table 8 above. While the initial inference time is 27.689 ms, each epoch of adaptation only introduce 3.292 ms (6.496 ms) additional computation time when combined with ERM (T3A), which is only 11.89% (23.46%) of the initial inference. This superior efficiency comes from (1) AdaRC only updating the hop-aggregation parameters and (2) the linear complexity of our PIC loss.
We also compare the computation time of AdaRC with other graph TTA algorithms. A significant disparity is observed: while the computation time for each step of adaptation in other graph TTA algorithms is several times that of inference, the adaptation time of our algorithm is merely 1/9 (1/4) of the inference time, making it almost negligible in comparison.
C.8 More architectures
Besides GPRGNN [7], our proposed AdaRC framework can also be integrated to more GNN architectures. We conduct experiments on Syn-cora dataset with three additional GNNs: APPNP [17], JKNet [43], and GCNII [6].
@@ C.8 More architectures
Figure 11: Performance of AdaRC on Syn-Cora with different GNN architectures.
The result are shown in Figure 11 above. Although different GNN architectures result in different performance on the target graph, AdaRC can consistently improve the accuracy. It shows that AdaRC is compatible with a wide range of GNN architectures.
@@ D.1 Datasets
We provide more details on the datasets used in the paper, including CSBM synthetic dataset and real-world datasets (Syn-Cora [51], Syn-Products [51], Twitch-E [31], and OGB-Arxiv [12]).
  • CSBM [8]. We use $N=5,000$ nodes on both source and target graph with $D=2,000$ features. Let ${\bm{\mu}}_{+}=\frac{0.03}{\sqrt{D}}\cdot\bm{1}_{D}$, ${\bm{\mu}}_{-}=-\frac{0.03}{\sqrt{D}}\cdot\bm{1}_{D}$, and $\Delta{\bm{\mu}}=\frac{0.02}{\sqrt{D}}\cdot\bm{1}_{D}$.
@@
  • Syn-Cora [51] and Syn-Products [51] are widely used datasets to evaluate model’s capability in handling homophly and heterophily. The Syn-Cora dataset is generated with various heterophily ratios based on modified preferential attachment process. Starting from an empty initial graph, new nodes are sequentially added into the graph to ensure the desired heterophily ratio. Node features are further generated by sampling node features from the corresponding class in the real-world Cora dataset. Syn-Products is generated in a similar way. For both dataset, we use $h=0.8$ as the source graph and $h=0.2$ as the target graph. We use non-overlapping train-test split over nodes on Syn-Cora to avoid label leakage.
  • Twitch-E [31] is a set of social networks, where nodes are Twitch users, and edges indicate friendships. Node attributes are the games liked, location and streaming habits of the user. We use ‘DE’ as the source graph and ‘ENGB’ as the target graph. We randomly drop a subset of homophily edges on the target graph to inject degree shift and homophily shift.
  • OGB-Arxiv [12] is a paper citation network of ARXIV papers, where nodes are ARXIV papers and edges are citations between these papers. Node attributes indicate the subject of each paper. We use a subgraph consisting of papers from 1950 to 2011 as the source graph, 2011 to 2014 as the validation graph, and 2014 to 2020 as the target graph. Similarly, we randomly drop a subset of homophily edges on the target graph to inject degree shift and homophily shift.
@@ E.1 Additional related works
Graph out-of-distribution generalization (graph OOD) aims to train a GNN model on the source graph that performs well on the target graph with unknown distribution shifts [18]. Existing graph OOD methods improve the model generalization by manipulating the source graph [29, 40], designing disentangled [24, 45] or casuality-based [19, 9] models, and exploiting various learning strategies [20, 52]. However, graph OOD methods focus on learning a universal model on source and target graphs, while not addressing model adaption to a specific target graph.
@@
Most GNN models follow the homophily assumption that neighboring nodes tend to share similar labels [16, 32]. Various message-passing [36, 50] and aggregation [7, 2, 51, 43] paradigms have been proposed to extend GNN models to heterophilic graphs. These GNN structures often embrace additional parameters, e.g., the aggregation weights for GPRGNN [7] and H2GCN [51], to handle both homophilic and heterophilic graphs. Such parameters provide the flexibility we need to adapt models to shifted graphs. However, these methods focus on the model design to handle either homophilic or heterophilic graph, without considering distribution shifts.
@@ E.2 Limitations
Since we mainly focus on the challenge of distribution shifts. Our proposed algorithm assumes that the source model should be able to learn class-clustered representations on the source graph, and should generalize well when there are no distribution shifts. In applications with extremely low signal-to-noise ratio, our algorithm’s improvement in accuracy might not be guaranteed. However, we would like to point out that this is a challenge faced by almost all TTA algorithms [49].
@@
Our proposed algorithm introduce additional computational overhead during testing. However, we quantify the additional computation time: it is minimal compared to the GNN inference time. Also, AdaRC is much more efficient that other graph TTA methods.

### v2

@@ 5 $\to$ 2 10 $\to$ 2 20 $\to$ 2 2 $\to$ 20 5 $\to$ 20 10 $\to$ 20
$+$ Matcha 88.78 $\pm$ 0.13 88.55 $\pm$ 0.44 88.10 $\pm$ 0.21 97.01 $\pm$ 1.00 97.24 $\pm$ 1.11 97.89 $\pm$ 0.25
C.4 Evolving target graphs
Our experiments in the main text mainly focused on adapting to a single target graph with a stable distribution. In this section, we test Matcha on a stream of target graphs with evolving structure shifts. Specifically, we used the Syn-Cora dataset, where the model was pre-trained on a source graph with a homophily of 0.8. It was then sequentially adapted to five target graphs with homophilies of 0.1, 0.7, 0.3, 0.9, and 0.2, simulating continuously changing homophily. The experimental setup, except for the sequence of target graphs, was identical to that in the Syn-Cora experiments described in the main text.
For example, in the column corresponding to the target graph with a homophily of 0.3, we compared two scenarios: (1) static: directly adapting from $h=0.8$ to $h=0.3$, and (2) evolving: sequentially adapting through $(h=0.8)\to(h=0.1)\to(h=0.7)\to(h=0.3)$.
Table 5: Accuracy (mean ± s.d. %) on Syn-Cora with static target graph and evolving target graph
Setting Method Target graph
$h=0.1$ $h=0.7$ $h=0.3$ $h=0.9$ $h=0.2$
Static ERM 63.19 $\pm$ 1.28 $8$$\pm$ 0.61 71.46 $\pm$ 0.62 97.15 $\pm$ 0.32 65.67 $\pm$ 0.35
Static Matcha 79.75 $\pm$ 1.04 90.57 $\pm$ 0.47 79.68 $\pm$ 0.73 97.40 $\pm$ 0.28 78.96 $\pm$ 1.08
Evolving Matcha 79.75 $\pm$ 1.04 90.65 $\pm$ 0.33 77.43 $\pm$ 0.62 97.31 $\pm$ 0.42 78.26 $\pm$ 1.02
The results, shown in Table 5 above, indicate that Matcha achieves performance on evolving graphs highly comparable to that on static graphs. This demonstrates Matcha’s ability to handle dynamic scenarios effectively, even under continuously changing graph structures.
C.5 Robustness to additional adversarial shift
While Matcha primarily targets natural structure shifts, inspired by (Jin et al., 2023), we test the robustness of Matcha against adversarial attacks by applying the PR-BCD attack (Geisler et al., 2021) on the target graph in our Syn-Cora experiments, varying the perturbation rate from 5% to 20%. The results are shown in Table 6. We found that while the accuracy of ERM dropped by 20.2%, the performance of Matcha only decreased by 2.3%. This suggests that our algorithm has some robustness to adversarial attacks, possibly due to the overlap between adversarial attacks and structure shifts. Specifically, we observed a decrease in homophily in the target graph under adversarial attack, indicating a similarity to structure shifts.
@@ C.5 Robustness to additional adversarial shift
C.6 Ablation study with different loss functions
We compare our proposed PIC loss with two existing surrogate losses: entropy (Wang et al., 2021) and pseudo-label (Liang et al., 2020). While PIC loss use the ratio form of $\sigma_{\text{intra}}^{2}$ and $\sigma_{\text{inter}}^{2}$, we also compare it with a difference form $\sigma_{\text{intra}}^{2}-\sigma_{\text{inter}}^{2}$, which also encourage larger $\sigma_{\text{inter}}^{2}$ and smaller $\sigma_{\text{intra}}^{2}$. The results are shown in Table 7: Our PIC loss has better performance under four structure shift scenarios.
@@ C.6 Ablation study with different loss functions
C.7 Hyperparameter sensitivity with different number of GPR steps $K$
Although the Matcha does not involve any hyperparameters other than the learning rate $\eta$ and number of adaptation rounds $T$, it may be combined with GNN models with different dimension of ${\bm{\gamma}}$. Therefore in this part, we combine Matcha with GPRGNN models using different $K$, i.e., the number of GPR steps, to test the robustness to different hyperparameter selection of the GNN model. Specifically, we tried values of $K$ ranging from 3 to 15 on Syn-Cora and Syn-Products datasets. Notice that in our experiments in 5.1, we use $K=9$. As shown in Table 8, Matcha remains effective under a wide range of $K$.
@@ C.7 Hyperparameter sensitivity with different number of GPR steps $K$
$+$ Matcha 72.09 $\pm$ 0.50 71.42 $\pm$ 0.65 70.58 $\pm$ 1.01 69.69 $\pm$ 1.06 69.48 $\pm$ 1.16 69.35 $\pm$ 0.66 69.72 $\pm$ 0.70
C.8 Computation time
Due to the need to adapt hop-aggregation parameters ${\bm{\gamma}}$, Matcha inevitably introduces additional computation costs, which vary depending on the chosen model, target graph, and base TTA algorithm. We documented the computation times for each component of ERM + Matcha and T3A + Matcha in our CSBM experiments:
@@ C.8 Computation time
    • Forward pass involves calculation of ${\bm{Z}}$ using the current ${\bm{\gamma}}$ and cached $\{{\bm{H}}^{(0)},\cdots,{\bm{H}}^{(K)}\}$, and prediction using the linear layer classifier (or with T3A algorithm). Since Matcha only updates ${\bm{\gamma}}$, $\{{\bm{H}}^{(0)},\cdots,{\bm{H}}^{(K)}\}$ can be cached without recomputation in each epoch. Note that other TTA algorithms could also adopt the same or similar caching strategies.
@@
ERM $+$ Matcha Adaptation (for each epoch) 3.292 $\pm$ 0.254 11.89%
   - Forward pass 1.224 $\pm$ 0.131 4.42%
   - Computing PIC loss 0.765 $\pm$ 0.019 2.76%
   - Back-propagation 1.189 $\pm$ 0.131 4.30%
   - Updating parameter 0.113 $\pm$ 0.001 0.41%
T3A $+$ Matcha Adaptation (for each epoch) 6.496 $\pm$ 0.333 23.46%
   - Forward pass 4.464 $\pm$ 0.248 16.12%
   - Computing PIC loss 0.743 $\pm$ 0.011 2.68%
   - Back-propagation 1.174 $\pm$ 0.167 4.24%
   - Updating parameter 0.115 $\pm$ 0.004 0.41%
We provide the computation time for each stage in Table 9 above. While the initial inference time is 27.689 ms, each epoch of adaptation only introduce 3.292 ms (6.496 ms) additional computation time when combined with ERM (T3A), which is only 11.89% (23.46%) of the initial inference. This superior efficiency comes from (1) Matcha only updating the hop-aggregation parameters and (2) the linear complexity of our PIC loss.
We also compare the computation time of Matcha with other graph TTA algorithms. A significant disparity is observed: while the computation time for each step of adaptation in other graph TTA algorithms is several times that of inference, the adaptation time of our algorithm is merely 1/9 (1/4) of the inference time, making it almost negligible in comparison.
C.9 Scalability
In the end of Section 4, we show the computational complexity of Matcha is linear to the number of nodes. To further validate this, we have conducted experiments on graphs of varying sizes from 1 million to 10 million nodes, and record the computation time for each epoch of adaptation. The results confirm that the computation time for Matcha indeed scales linearly with graph size, demonstrating its efficiency even for very large graphs.
Figure 12: Scalability of Matcha
C.10 More architectures
Besides GPRGNN (Chien et al., 2021), our proposed Matcha framework can also be integrated to more GNN architectures. We conduct experiments on Syn-cora dataset with three additional GNNs: APPNP (Klicpera et al., 2019), JKNet (Xu et al., 2018), and GCNII (Chen et al., 2020).
@@ C.10 More architectures
Notice that Tent can only be applied to models with batch normalization layers, which are not included for JKNet in GCNII in our implementation.
Table 10: Accuracy (mean $\pm$ s.d.) on Syn-Cora with different GNN architectures
Method GPRGNN APPNP JKNet GCNII
ERM 65.67 $\pm$ 0.35 70.24 $\pm$ 0.88 47.87 $\pm$ 0.90 67.95 $\pm$ 1.33
$+$ Matcha 78.96 $\pm$ 1.08 80.63 $\pm$ 0.35 51.57 $\pm$ 2.09 74.33 $\pm$ 0.45
T3A 68.25 $\pm$ 1.10 70.98 $\pm$ 0.86 47.93 $\pm$ 0.85 68.20 $\pm$ 1.31
$+$ Matcha 78.40 $\pm$ 1.04 80.70 $\pm$ 0.38 51.84 $\pm$ 1.87 74.96 $\pm$ 0.23
Tent 66.26 $\pm$ 0.38 70.15 $\pm$ 1.08 - -
$+$ Matcha 78.87 $\pm$ 1.07 80.72 $\pm$ 0.18 - -
AdaNPC 67.34 $\pm$ 0.76 70.53 $\pm$ 0.76 47.93 $\pm$ 0.77 68.39 $\pm$ 1.18
$+$ Matcha 77.45 $\pm$ 0.62 80.11 $\pm$ 0.61 48.32 $\pm$ 0.69 74.44 $\pm$ 0.35
GTrans 68.60 $\pm$ 0.32 73.50 $\pm$ 0.62 51.38 $\pm$ 0.58 74.08 $\pm$ 1.26
$+$ Matcha 83.49 $\pm$ 0.78 85.17 $\pm$ 0.43 53.76 $\pm$ 2.26 80.50 $\pm$ 0.40
SOGA 67.16 $\pm$ 0.72 78.62 $\pm$ 0.48 47.96 $\pm$ 0.55 66.87 $\pm$ 1.50
$+$ Matcha 79.03 $\pm$ 1.10 80.88 $\pm$ 0.56 52.05 $\pm$ 1.64 74.39 $\pm$ 0.29
GraphPatcher 63.01 $\pm$ 2.29 57.49 $\pm$ 1.83 45.38 $\pm$ 1.00 67.05 $\pm$ 1.54
$+$ Matcha 80.99 $\pm$ 0.50 81.38 $\pm$ 0.88 46.78 $\pm$ 1.71 74.46 $\pm$ 0.50
The result are shown in Table 10 above. Although different GNN architectures result in different performance on the target graph, Matcha can consistently improve the accuracy. It shows that Matcha is compatible with a wide range of GNN architectures.
@@ D.1 Datasets
We provide more details on the datasets used in the paper, including CSBM synthetic dataset and real-world datasets (Syn-Cora (Zhu et al., 2020), Syn-Products (Zhu et al., 2020), Twitch-E (Rozemberczki et al., 2021), and OGB-Arxiv (Hu et al., 2020)).
  • CSBM (Deshpande et al., 2018). We use $N=5,000$ nodes on both source and target graph with $D=2,000$ features. Let ${\bm{\mu}}_{+}=\frac{0.03}{\sqrt{D}}\cdot\bm{1}_{D}$, ${\bm{\mu}}_{-}=-\frac{0.03}{\sqrt{D}}\cdot\bm{1}_{D}$, and $\Delta{\bm{\mu}}=\frac{0.02}{\sqrt{D}}\cdot\bm{1}_{D}$.
@@
  • Syn-Cora (Zhu et al., 2020) and Syn-Products (Zhu et al., 2020) are widely used datasets to evaluate model’s capability in handling homophly and heterophily. The Syn-Cora dataset is generated with various heterophily ratios based on modified preferential attachment process. Starting from an empty initial graph, new nodes are sequentially added into the graph to ensure the desired heterophily ratio. Node features are further generated by sampling node features from the corresponding class in the real-world Cora dataset. Syn-Products is generated in a similar way. For both dataset, we use $h=0.8$ as the source graph and $h=0.2$ as the target graph. We use non-overlapping train-test split over nodes on Syn-Cora to avoid label leakage.
  • Twitch-E (Rozemberczki et al., 2021) is a set of social networks, where nodes are Twitch users, and edges indicate friendships. Node attributes are the games liked, location and streaming habits of the user. We use ‘DE’ as the source graph and ‘ENGB’ as the target graph. We randomly drop a subset of homophily edges on the target graph to inject degree shift and homophily shift.
  • OGB-Arxiv (Hu et al., 2020) is a paper citation network of ARXIV papers, where nodes are ARXIV papers and edges are citations between these papers. Node attributes indicate the subject of each paper. We use a subgraph consisting of papers from 1950 to 2011 as the source graph, 2011 to 2014 as the validation graph, and 2014 to 2020 as the target graph. Similarly, we randomly drop a subset of homophily edges on the target graph to inject degree shift and homophily shift.
@@ E.1 Additional related works
Graph neural networks (GNNs) have shown great success in various graph applications (Kipf & Welling, 2017; Velickovic et al., 2017; Rozemberczki et al., 2021; Qiu et al., 2022; 2023; Hu et al., 2020; Wang et al., 2023a; Pareja et al., 2020). In Section 2, we briefly introduced selected related works in test-time adaptation and graph domain adaptation. Here with discuss more related works in test-time adaptation, graph out-of-distribution generalization, and homophily-adaptive GNN models.
More test-time adaptation
Another important category of TTA algorithms, beyond those discussed in the main text, leverages data augmentation (Ding et al., 2022; Fu et al., 2023). These methods apply multiple augmentations to the input and enhance prediction robustness through techniques such as ensembling (Hendrycks et al., 2020) or minimizing marginal entropy (Zhang et al., 2022). In the context of graphs, augmentation can also be utilized for test-time adaptation (Xu et al., 2022; Liu et al., 2024b; Zeng et al., 2024). For instance, GTrans (Jin et al., 2023) and DropEdge (Rong et al., 2020) employ augmentation strategies for this purpose. GTrans adopts DropEdge as an augmentation technique and incorporates a contrastive loss that maximizes the similarity between original nodes and their augmented views while penalizing their similarity to negative samples. However, augmentation-based approaches inevitably introduce significant additional computational costs.
Graph out-of-distribution generalization (graph OOD) aims to train a GNN model on the source graph that performs well on the target graph with unknown distribution shifts (Li et al., 2022a). Existing graph OOD methods improve the model generalization by manipulating the source graph (Park et al., 2021; Wu et al., 2022), designing disentangled (Ma et al., 2019; Yang et al., 2020; Yan et al., 2024c; Zeng et al., 2023) or casuality-based (Li et al., 2023; Fan et al., 2024) models, and exploiting various learning strategies (Li et al., 2022b; Zhu et al., 2021b). However, graph OOD methods focus on learning a universal model on source and target graphs, while not addressing model adaption to a specific target graph.
@@
Most GNN models follow the homophily assumption that neighboring nodes tend to share similar labels (Kipf & Welling, 2017; Velickovic et al., 2017; Yan et al., 2024a). Various message-passing (Wang & Zhang, 2022; Zhu et al., 2021a; Xu et al., 2024; Yan et al., 2024b) and aggregation (Chien et al., 2021; Bo et al., 2021; Zhu et al., 2020; Xu et al., 2018; 2023) paradigms have been proposed to extend GNN models to heterophilic graphs. These GNN structures often embrace additional parameters, e.g., the aggregation weights for GPRGNN (Chien et al., 2021) and H2GCN (Zhu et al., 2020), to handle both homophilic and heterophilic graphs. Such parameters provide the flexibility we need to adapt models to shifted graphs. However, these methods focus on the model design to handle either homophilic or heterophilic graph, without considering distribution shifts.
@@ E.2 Limitations
Since we mainly focus on the challenge of distribution shifts. Our proposed algorithm assumes that the source model should be able to learn class-clustered representations on the source graph, and should generalize well when there are no distribution shifts. In applications with extremely low signal-to-noise ratio, our algorithm’s improvement in accuracy might not be guaranteed. However, we would like to point out that this is a challenge faced by almost all TTA algorithms (Zhao et al., 2023).
@@
Our proposed algorithm introduce additional computational overhead during testing. However, we quantify the additional computation time: it is minimal compared to the GNN inference time. Also, Matcha is much more efficient that other graph TTA methods.